you are walking from your physics class at a speed of 1.0 m/s. your friend leaves two minutes later at the rate of 1.3 m/s. how long will it take you friend to catch up to you?
Answer by richwmiller(17219)


Gaining speed 1.3 -1= 0.3 m/s
Headstart 1*120 =120 meters
Time to catch up 120/0.3=400.0 sec
Algebraic solution
1(t+120)=1.3t
1t+120=1.3t
120=1.3t-1t
120=0.3t
t=400.0 sec
t=400/60=6 2/3 minutes
Distance to meeting 400.0*1.3=520 meters